hon flies from atlanta, georgia to san francisco,california. It take 5.6 hours to travel 2100 miles against the head wind. at the same time, Debby flies fro msan fransico to atlanta her plane travels with the same average airspeed , but with a tail wind, but her flight only takes 4.8 hours. Write a system of equations that relates time,airspeed, and wind speed to distance for each traveler.Solve the system to find the airspeed
Answer by jorel1380(2518)

2100/4.8=437.5
2100/5.6=375
437.5-375/2=31.25
437.5-31.25=406.25
375+31.25=406.25
The speed of the plane in still air is 406.25 mph. The wind speed is 31.25 mph..